OpenStudy (anonymous):

In a circle, a 60 degree sector has area 25 pi ft^2. What is the circumference of the circle? Leave your answer in terms of pi and simplest radical form. I don't have any idea where to start.. Help!

OpenStudy (anonymous):

area of a sector = (C/360)*r^2 where C is the angle of the sector in degrees circumference = 2pi*r so solve for r in the top equation, then plug in r into the second equation, ya? ^_^

OpenStudy (anonymous):

here is the first step.. 25*pi = (60/360)*r^2

OpenStudy (anonymous):

So R = 2.04, right?

OpenStudy (anonymous):

Or 2.04 pi? Sorry! I'm not good with math..

OpenStudy (anonymous):

\[\frac {25*\pi} { \frac{60}{360}} = r^2\] \[\frac {25*\pi} { \frac{60}{360}} = r^2\

OpenStudy (anonymous):

\[\frac {360*25*\pi} {60} = r^2\]

OpenStudy (anonymous):

can you work it out from there?